OpenStudy (anonymous):

Find AB giving the following cordnites A(0,0) and B(0,-6) (distance formula)

OpenStudy (anonymous):

Use \(\sqrt{\Delta x^2+\Delta y^2}.\)

OpenStudy (anonymous):

would the answer be 12?

OpenStudy (anonymous):

or 6 ? @Limitless

OpenStudy (anonymous):

It's \(6\). You have \(\sqrt{\Delta x^2+\Delta y^2}=\sqrt{(0-0)^2+(0-(-6))^2}=\sqrt{0^2+6^2}=\sqrt{6^2}=6.\) Does this make sense?

OpenStudy (anonymous):

so am i right ?

OpenStudy (anonymous):

What answer do you think it is?

OpenStudy (anonymous):

6???

OpenStudy (anonymous):

its 36 huh @Limitless

OpenStudy (anonymous):

6 is correct.

OpenStudy (anonymous):

Read my little line of math there and it's clear why this is so.

OpenStudy (anonymous):

lol.okay thanls